OpenStudy (anonymous):

Can someone help me simplify this complex fraction? 10/x - 5 / 5

OpenStudy (arlen):

10/x-1 (10-x)/x

hartnn (hartnn):

yes,many can help u.

OpenStudy (anonymous):

5/5 =1

OpenStudy (anonymous):

so would itbe 1/x?

OpenStudy (arlen):

\[\frac{ 10 }{ x }-1=\frac{ 10-x }{ x }\]

OpenStudy (anonymous):

@arlen Directly did it

OpenStudy (anonymous):

ANd after that he took L.C.M

OpenStudy (anonymous):

So is @arlen saying that the final fraction is 10-x/x?

OpenStudy (arlen):

yes kak

OpenStudy (anonymous):

the multiple choices are 1/x, 1, and 2.. would it be 1/x ?

hartnn (hartnn):

can u draw your question kak ??

OpenStudy (anonymous):

or 2-x/ x

OpenStudy (anonymous):



hartnn (hartnn):

ohh

OpenStudy (arlen):

10-5x/5= 2-x

OpenStudy (anonymous):

2-x/ x ! Thank you so much :)

OpenStudy (arlen):

yes....sorry i forgot denominator

OpenStudy (anonymous):

yes, thats okay thank you so much! Iw ill be posting more complex fraction problems haha

OpenStudy (arlen):

it's cool kak

hartnn (hartnn):

make sure u post it correctly or draw it.

OpenStudy (anonymous):

okay thank you!